OpenStudy (anonymous):

Mary and Marsha have a combined age of 36. Marsha is twice as old as Mary. How old is Mary?

Mehek (mehek14):

2x + x = 36

Mehek (mehek14):

2x + x = 3x

OpenStudy (thatonegirl_):

well start with what you know. Make Mary=x. Now you can have marsha=2x (twice the age of mary). Then set x+2x=36 (because it is the combined age of both of them). Solve for x and that gives you the age of Mary. If you wanted to know Marsha's, just double it.

Mehek (mehek14):

3x = 36 divide 36/3 @Caitlyn_Morgan

OpenStudy (anonymous):

Thank you both! ^-^

Mehek (mehek14):

what did you get?

OpenStudy (anonymous):

I got 12

Mehek (mehek14):

ok correct

OpenStudy (thatonegirl_):

Ur welcome :)